# On the Mode-Seeking Properties of Langevin Dynamics

### v1

@@ A.3 Proof of Theorem 3: Langevin Dynamics under Sub-Gaussian Mixtures
A.4 Proof of Theorem 4: Annealed Langevin Dynamics under Sub-Gaussian Mixtures
@@
A.5 Proof of Theorem 5: Convergence Analysis of Chained Langevin Dynamics
@@
Thus we finish the proof of Theorem 5.
Appendix B Additional Experiments
Algorithm Setup: Our choices of algorithm hyperparameters are based on [2]. We consider $L=10$ different standard deviations such that $\left\{{\lambda_{i}}\right\}_{i\in[L]}$ is a geometric sequence with $\lambda_{1}=1$ and $\lambda_{10}=0.01$. For annealed Langevin dynamics with $T$ iterations, we choose the noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[T]}$ by repeating every element of $\left\{{\lambda_{i}}\right\}_{i\in[L]}$ for $T/L$ times and we set the step size as $\delta_{t}=2\times 10^{-5}\cdot{\sigma}_{t}^{2}/{\sigma}_{T}^{2}$ for every $t\in[T]$. For vanilla Langevin dynamics with $T$ iterations, we use the same step size as annealed Langevin dynamics. For chained Langevin dynamics with $T$ iterations, the patch size $Q$ is chosen depending on different tasks. For every patch of chained Langevin dynamics, we choose the noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]}$ by repeating every element of $\left\{{\lambda_{i}}\right\}_{i\in[L]}$ for $TQ/dL$ times and we set the step size as $\delta_{t}=2\times 10^{-5}\cdot{\sigma}_{t}^{2}/{\sigma}_{TQ/d}^{2}$ for every $t\in[TQ/d]$.
B.1 Synthetic Gaussian Mixture Model
@@
Since the distribution is given, we assume that the sampling algorithms have access to the ground-truth score function. We set the batch size as 1000 and patch size $Q=10$ for chained Langevin dynamics. We use $T\in\left\{{10^{3},10^{4},10^{5},10^{6}}\right\}$ iterations for vanilla, annealed, and chained Langevin dynamics. The initial samples are i.i.d. chosen from $P^{(0)}$, $P^{(1)}$, or $P^{(2)}$, and the results are presented in Figures 1, 4, and 5 respectively. The two subfigures above the dashed line illustrate the samples from the initial distribution and target distribution, and the subfigures below the dashed line are the samples generated by different algorithms. A sample ${\mathbf{x}}$ is clustered in mode 1 if it satisfies $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}\leq 5d$ and $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}\leq\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}$; in mode 2 if $\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}\leq 5d$ and $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}>\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}$; and in mode 0 otherwise. The experiments were run on an Intel Xeon CPU with 2.90GHz.
B.2 Image Datasets
@@
Our implementation and hyperparameter selection are based on [2]. During training, we i.i.d. randomly flip an image with probability 0.5 to construct the two modes (i.e., original and flipped images). All models are optimized by Adam with learning rate 0.001 and batch size 128 for a total of 200000 training steps, and we use the model at the last iteration to generate the samples. We perform experiments on MNIST [45] (CC BY-SA 3.0 License) and Fashion-MNIST [46] (MIT License) datasets and we set the patch size as $Q=14$.
@@
For the score networks of vanilla and annealed Langevin dynamics, following from [2], we use the 4-cascaded RefineNet [49], a modern variant of U-Net [47] with residual design. For the score networks of chained Langevin dynamics, we use the official PyTorch implementation of an LSTM network [50] followed by a linear layer. For MNIST and Fashion-MNIST datasets, we set the input size of the LSTM as $Q=14$, the number of features in the hidden state as 1024, and the number of recurrent layers as 2. The inputs of LSTM include inputting tensor, hidden state, and cell state, and the outputs of LSTM include the next hidden state and cell state, which can be fed to the next input. To estimate the noisy score function, we first input the noise level ${\sigma}$ (repeated for $Q$ times to match the input size of LSTM) and all-0 hidden and cell states to obtain an initialization of the hidden and cell states. Then, we divide a sample into $d/Q$ patches and input the sequence of patches to the LSTM. For every output hidden state corresponding to one patch, we apply a linear layer of size $1024\times Q$ to estimate the noisy score function of the patch.
@@
To generate samples, we use $T\in\left\{{3000,10000,30000,100000}\right\}$ iterations for vanilla, annealed, and chained Langevin dynamics. The initial samples are chosen as either original or flipped images from the dataset, and the results for MNIST and Fashion-MNIST datasets are presented in Figures 2, 6, 3, and 7 respectively. The two subfigures above the dashed line illustrate the samples from the initial distribution and target distribution, and the subfigures below the dashed line are the samples generated by different algorithms. High-quality figures generated by annealed and chained Langevin dynamics for $T=100000$ iterations are presented in Figures 8 and 9.
@@
Figure 4: Samples from a mixture of three Gaussian modes generated by vanilla, annealed, and chained Langevin dynamics. Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 1.
Figure 5: Samples from a mixture of three Gaussian modes generated by vanilla, annealed, and chained Langevin dynamics. Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 2.
Figure 6: Samples from a mixture distribution of the original and flipped images from the MNIST dataset generated by vanilla, annealed, and chained Langevin dynamics. The samples are initialized as flipped images from MNIST.
Figure 7: Samples from a mixture distribution of the original and flipped images from the Fashion-MNIST dataset generated by vanilla, annealed, and chained Langevin dynamics. The samples are initialized as flipped images from Fashion-MNIST.
Figure 8: Samples from a mixture distribution of the original and flipped images from the MNIST dataset generated by annealed and chained Langevin dynamics for $T=100000$ iterations. The samples are initialized as the original or flipped images from MNIST.
Figure 9: Samples from a mixture distribution of the original and flipped images from the Fashion-MNIST dataset generated by annealed and chained Langevin dynamics for $T=100000$ iterations. The samples are initialized as the original or flipped images from Fashion-MNIST.

### v2

@@ A.3 Proof of Theorem 2: Langevin Dynamics under Sub-Gaussian Mixtures
Appendix B Theoretical Analysis on Annealed Langevin Dynamics
B.1 Annealed Langevin Dynamics in Gaussian Mixtures
In the following Theorem 3 we extend the result in Theorem 1 to annealed Langevin dynamics with bounded noise levels.
Theorem 3.
Proof of Theorem 3.
From equation 2 we note that the perturbed distribution is the convolution of the original distribution and a Gaussian random variable, i.e., for random variables ${\mathbf{z}}\sim p$ and ${\mathbf{t}}\sim{\mathcal{N}}(\bm{0}_{d},{\bm{I}}_{d})$, their sum ${\mathbf{z}}+{\mathbf{t}}\sim p_{\sigma}$ follows the perturbed distribution with noise level ${\sigma}$. Therefore, a perturbed (sub)Gaussian distribution remains (sub)Gaussian. We formalize this property in Proposition 5.
Proposition 5.
Proof of Proposition 5.
By the definition in equation 2, we have
Similar to the proof of Theorem 1, we decompose
Proposition 6.
Proof of Proposition 6.
    we have that for all $t\in[T]$,
    By Lemma 8 we have that with probability at least $1-\exp(-\Omega(d))$,
    Combining the two inequalities implies the desired bound
Therefore we complete the proof of Proposition 6. ∎
Finally, combining equation 20 and Proposition 6 finishes the proof of Theorem 3. ∎
B.2 Annealed Langevin Dynamics in Sub-Gaussian Mixtures
Finally, we slightly modify Assumption 2 and extend our results to annealed Langevin dynamics (with bounded noise levels) under sub-Gaussian mixtures in Theorem 4.
@@ B.2 Annealed Langevin Dynamics in Sub-Gaussian Mixtures
Finally, combining equation 21 and Proposition 7 finishes the proof of Theorem 4. ∎
Appendix C Convergence Analysis of Chained Langevin Dynamics
Proof of Proposition 1.
@@ Appendix C Convergence Analysis of Chained Langevin Dynamics
Thus we finish the proof of Proposition 1. ∎
Appendix D Experimental Details and Additional Experiments
Algorithm Setup: Our choices of algorithm hyperparameters are based on [4] and [26]. For $\sigma_{\max}=1$, following from [4], we consider $L=10$ different standard deviations such that $\left\{{\lambda_{i}}\right\}_{i\in[L]}$ is a geometric sequence with $\lambda_{1}=1$ and $\lambda_{10}=0.01$. For annealed Langevin dynamics with $T$ iterations, we choose the noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[T]}$ by repeating every element of $\left\{{\lambda_{i}}\right\}_{i\in[L]}$ for $T/L$ times and we set the step size as $\delta_{t}=2\times 10^{-5}\cdot{\sigma}_{t}^{2}/{\sigma}_{T}^{2}$ for every $t\in[T]$. For vanilla Langevin dynamics with $T$ iterations, we use the same step size as annealed Langevin dynamics. For Chained-VLD and Chained-ALD, the patch size $Q$ is chosen depending on different tasks. For every patch of Chained-ALD, we choose the noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]}$ by repeating every element of $\left\{{\lambda_{i}}\right\}_{i\in[L]}$ for $TQ/dL$ times and we set the step size as $\delta_{t}=2\times 10^{-5}\cdot{\sigma}_{t}^{2}/{\sigma}_{TQ/d}^{2}$ for every $t\in[TQ/d]$. The step size of Chained-VLD is the same as Chained-ALD.
We would like to highlight that the inference time of Chained-LD is significantly lower than vanilla LD in practice. Our theoretical comparison between Chained-LD and vanilla LD is based on iteration complexity, i.e., the number of queries to the score function $\nabla\log P(x^{(q)}|x^{(1)},\cdots,x^{(q-1)})$ or $\nabla\log P(x)$. Since Chained-LD only updates one patch at every iteration while vanilla LD updates the whole image, Chained-LD will be significantly faster than vanilla LD.
D.1 Synthetic Gaussian Mixture Model
@@
Since the distribution is given, we assume that the sampling algorithms have access to the ground-truth score function. We set the batch size as 1000 and patch size $Q=10$ for chained Langevin dynamics. We use $T\in\left\{{10^{3},10^{4},10^{5},10^{6}}\right\}$ iterations for vanilla and chained Langevin dynamics. A sample ${\mathbf{x}}$ is clustered in mode 1 if it satisfies $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}\leq 5d$ and $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}\leq\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}$; in mode 2 if $\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}\leq 5d$ and $\left\|{{\mathbf{x}}-{\bm{\mu}}_{1}}\right\|^{2}>\left\|{{\mathbf{x}}-{\bm{\mu}}_{2}}\right\|^{2}$; and in mode 0 otherwise. The initial samples are i.i.d. chosen from $P^{(0)}$, $P^{(1)}$, or $P^{(2)}$, and the results are presented in Figures 2, 5, and 6 respectively. The two subfigures above the dashed line illustrate the samples from the initial distribution and target distribution, and the subfigures below the dashed line are the samples generated by different algorithms. Furthermore, in Figures 7, 8 and 9 we demonstrate the effect of different values of $Q\in\{1,4,10,20,50\}$ on the mode-seeking tendencies of Chained-LD. We can observe that for dimension $d=100$, a moderate patch size $Q\in\{1,4,10\}$ has similar performance, a large patch size $Q=20$ needs more steps to find the other two modes, while an overly-large patch size $Q=50$ almost cannot find other modes.
D.2 Score Function Estimator
In realistic scenarios, since we do not have direct access to the (perturbed) score function, [4] proposed the Noise Conditional Score Network (NCSN) ${\mathbf{s}}_{\bm{\theta}}({\mathbf{x}},{\sigma})$ to jointly estimate the scores of all perturbed data distributions, i.e.,
Assuming the NCSN has enough capacity and sufficient training samples, ${\mathbf{s}}_{\bm{\theta}^{*}}({\mathbf{x}},{\sigma})$ minimizes the loss ${\mathcal{L}}\left(\bm{\theta};\left\{{{\sigma}_{t}}\right\}_{t\in[T]}\right)$ if and only if ${\mathbf{s}}_{\bm{\theta}^{*}}({\mathbf{x}},{\sigma}_{t})=\nabla_{\mathbf{x}}\log P_{{\sigma}_{t}}({\mathbf{x}})$ almost surely for all $t\in[T]$.
In Chained Langevin dynamics, an ideal conditional score function estimator ${\mathbf{s}}_{\bm{\theta}}$ could jointly estimate the scores of all perturbed conditional patch distribution, i.e., $\forall{\sigma}\in\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]},q\in[d/Q]$,
Then, combining the objectives gives the following loss
As shown in [44], an estimator ${\mathbf{s}}_{\bm{\theta}}$ with enough capacity and sufficient training samples minimizes the loss ${\mathcal{L}}$ if and only if ${\mathbf{s}}_{\bm{\theta}}$ outputs the scores of all perturbed conditional patch distribution almost surely.
D.3 Image Datasets
@@
Our implementation and hyperparameter selection are based on [4] and [26]. During training, we i.i.d. randomly flip an image with probability 0.5 to construct the two modes (i.e., original and flipped images). All models are optimized by Adam with learning rate 0.001 and batch size 128 for a total of 200000 training steps, and we use the model at the last iteration to generate the samples. We perform experiments on MNIST [41] (CC BY-SA 3.0 License) and Fashion-MNIST [42] (MIT License) datasets and we set the patch size as $Q=14$.
@@
For the score networks of chained annealed Langevin dynamics (Chained-ALD), we use the official PyTorch implementation of an LSTM network [45] followed by a linear layer. For MNIST and Fashion-MNIST datasets, we set the input size of the LSTM as $Q=14$, the number of features in the hidden state as 1024, and the number of recurrent layers as 2. The inputs of LSTM include inputting tensor, hidden state, and cell state, and the outputs of LSTM include the next hidden state and cell state, which can be fed to the next input. To estimate the noisy score function, we first input the noise level ${\sigma}$ (repeated for $Q$ times to match the input size of LSTM) and all-0 hidden and cell states to obtain an initialization of the hidden and cell states. Then, we divide a sample into $d/Q$ patches and input the sequence of patches to the LSTM. For every output hidden state corresponding to one patch, we apply a linear layer of size $1024\times Q$ to estimate the noisy score function of the patch.
@@
To generate samples, we use $T\in\left\{{10000,30000,100000}\right\}$ iterations for annealed Langevin dynamics (ALD) and Chained-ALD. The initial samples are chosen as either original or flipped images from the dataset, and the results for MNIST and Fashion-MNIST datasets are presented in Figures 11, 3, 4, and 12 respectively. The two subfigures above the dashed line illustrate the samples from the initial distribution and target distribution, and the subfigures below the dashed line are the samples generated by different algorithms.
@@
Figure 5: Samples from a mixture of three Gaussian modes generated by vanilla Langevin dynamics (VLD) and chained vanilla Langevin dynamics (Chained-VLD). Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 1.
Figure 6: Samples from a mixture of three Gaussian modes generated by vanilla Langevin dynamics (VLD) and chained vanilla Langevin dynamics (Chained-VLD). Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 2.
Figure 7: Samples from a mixture of three Gaussian modes generated by vanilla Langevin dynamics (VLD) and chained vanilla Langevin dynamics (Chained-VLD) with patch size $Q\in\{1,4,10,20,50\}$. Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 0.
Figure 8: Samples from a mixture of three Gaussian modes generated by vanilla Langevin dynamics (VLD) and chained vanilla Langevin dynamics (Chained-VLD) with patch size $Q\in\{1,4,10,20,50\}$. Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 1.
Figure 9: Samples from a mixture of three Gaussian modes generated by vanilla Langevin dynamics (VLD) and chained vanilla Langevin dynamics (Chained-VLD) with patch size $Q\in\{1,4,10,20,50\}$. Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 2.
Figure 10: Samples from a mixture of three Gaussian modes generated by Langevin dynamics and chained Langevin dynamics with patch size $Q=10$. Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 0.
Figure 11: Samples from a mixture distribution of the original and flipped images from the MNIST dataset generated by annealed Langevin dynamics (ALD) and chained annealed Langevin dynamics (Chained-ALD) for different numbers of iterations. The maximum noise level $\sigma_{\max}$ is set to be 1 or 50. The samples are initialized as original images from MNIST.
Figure 12: Samples from a mixture distribution of the original and flipped images from the Fashion-MNIST dataset generated by annealed Langevin dynamics (ALD) and chained annealed Langevin dynamics (Chained-ALD) for different numbers of iterations. The maximum noise level $\sigma_{\max}$ is set to be 1 or 50. The samples are initialized as flipped images from FashionMNIST.